OpenStudy (anonymous):

) An objectis thrown upward from the edge of a tall building with a velocity of 10m/s. Where will the object be 3s after it is thrown? Take g = 10m/s^2

OpenStudy (alfie):

Use the uniform acceleration formula with a= 10m/s^2 and v0 = 10m/s

OpenStudy (turingtest):

\[s(t)=-g t^2/2+v_0t+x_0\]\[s(3)=-10(3^2/2)+10(3)=-15\]or 15m downward from its initial point

OpenStudy (turingtest):

remember if up is positive a=-10m/s